OpenStudy (anonymous):

show that d(tan^3x) / dx = 3sec^4x - 3sec^2x please help,

OpenStudy (anonymous):

d(tan^3x) / dx = 3tan^2x*sec^2x tan^2x=sec^2x-1 (3sec^2x-3)*sec^2x=3sec^4x-3sec^2x

OpenStudy (yttrium):

3tan^2(x) sec^2(x) = 3sec^4(x) -3sec^2 (x) Using pythagorean identities, 3(sec^2 (x) - 1) sec^2(x) = 3sec^4(x) -3sec^2 (x) Therefore, 3sec^4(x) - 3sec^2(x) = 3sec^4(x) -3sec^2 (x) Got it?